OpenStudy (anonymous):

Find the length of the curve y=(x^2/2) - (lnx/2) ; 2 less than or equal to x which is less than or equal to 4.

OpenStudy (anonymous):

parameterize the curve as (t, (t^2/2 - ((ln t)/2)) arc length is \[\int\limits_{t_{1}}^{t_{2}} ds\]

OpenStudy (anonymous):

where \[ds = \sqrt { (dx/dt)^2 + (dy/dt)^2} dt\]

OpenStudy (anonymous):

u understand>??

OpenStudy (anonymous):

hmmm i think so, i'm trying to follow along

OpenStudy (anonymous):

so whts dx/dt and dy/dt?

OpenStudy (anonymous):

derivatives of "x^2" and "ln(x)"?

OpenStudy (anonymous):

ill mae it easy x=t y=t^2 / 2 - (lnx)/2 now whts dx/dt and dy/dt?

OpenStudy (anonymous):

sorry, but i still don't get it :/

OpenStudy (anonymous):

DO U KNOW PARAMETERIZATION?

OpenStudy (anonymous):

ok lets try something else \[ds = \sqrt{ 1 + (dy/dx)^2 } dx\]

OpenStudy (anonymous):

now can u tell me dy/dx??

OpenStudy (anonymous):

u ther?

OpenStudy (anonymous):

no i don't know paramertization, sorry

OpenStudy (anonymous):

ok tell me dy/dx now

OpenStudy (anonymous):

where r u?

OpenStudy (anonymous):

i'm right here. i'm just looking up parameterization

OpenStudy (anonymous):

dont

OpenStudy (anonymous):

u dont need to

OpenStudy (anonymous):

just tell me the value of dy/dx?

OpenStudy (anonymous):

okay umm would it be t- 1/x?

OpenStudy (anonymous):

y = (x^2 - ln x)/2 dy/dx = x - 1/2x

OpenStudy (anonymous):

how did you get x? oh wait no, it's be cause t=x nvm

OpenStudy (anonymous):

leave t out now

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so now i told u ds = sqrt [ 1 + (dy/dx)^2] dx

OpenStudy (anonymous):

now give me ds..whats ds? in terms of x?

OpenStudy (anonymous):

okay hang on

OpenStudy (anonymous):

would ds = (x^2 - 2x +2)/(4x^2) ?

OpenStudy (anonymous):

yes most prob

OpenStudy (anonymous):

so now evaluate this wrt integral dx from x1 to x2

OpenStudy (anonymous):

square root of this

OpenStudy (anonymous):

its not right ...ds

OpenStudy (anonymous):

\[ds = \sqrt { x^2 + 1/4x^2}dx\]

OpenStudy (anonymous):

oooohhhh okay i'm writing this down and taking notes